OpenStudy (anonymous):

What is the equation of the line expressed in slope-intercept form? MEDAL FOR WHOMEVER HELPS....plz

OpenStudy (anonymous):

here is the graph... not sure if the other is opening or not...

OpenStudy (amilapsn):

what would be the slope?

OpenStudy (anonymous):

y = –3x – 1 y = 3x + 2 y = 2x + 3 y = –2x + 3 these are the answer choices

OpenStudy (amilapsn):

Slope is positive right?

OpenStudy (anonymous):

y=mx+b in example... y=7x+2 7 would be the slope and 2 would be the y-intercept

OpenStudy (amilapsn):

yes..

OpenStudy (anonymous):

i am not good at doing it from a graph tho:/

OpenStudy (amilapsn):

slope is just y change to x change

OpenStudy (amilapsn):

k?

OpenStudy (anonymous):

i think the slope is negative so it would have to be either the first choice or the last by process of elimination

OpenStudy (amilapsn):

just try to climb the graph from left to right.... do you have to climb up or down..?

OpenStudy (anonymous):

it has to be D or the last choice because the y-intercept must be positive based on the graph....i think...

OpenStudy (anonymous):

so the answer is y = –2x + 3

OpenStudy (amilapsn):

No!!! \(\color{red}{\Large\sf{STOP!!!}}\)

OpenStudy (anonymous):

???

OpenStudy (amilapsn):

just tell me whether the graph is ascending or descending.....

OpenStudy (anonymous):

jpg of the graph is attached

OpenStudy (amilapsn):

ascending: \(\Large\sf{y}\) is increasing when \(\Large\sf{x}\) is increasing descending:\(\Large\sf{y}\) is decreasing when \(\Large\sf{x}\) is increasing

OpenStudy (anonymous):

y is in quadrant 2 and x is in quadrant 3 y is in positive quadrant and x is in negative quadrant

OpenStudy (anonymous):

that did not help bc the answer it gave does not fit with the answer choices you have to choose from